OpenStudy (anonymous):

solve the following system of equations x + 6y = 6 (1) x = 7 - 6y (2)

OpenStudy (anonymous):

no solution

OpenStudy (mathteacher1729):

Why is that the correct answer, Adrianosst?

OpenStudy (anonymous):

because x=6-6y and x=7-6y. So 6=7?

OpenStudy (mathteacher1729):

That's a little clearer, but how did you obtain 6 = 7 ? You did something with x = 6-6y and x = 7 - 6y I guess... :)

OpenStudy (anonymous):

i brought them both to the x=blah blah form

OpenStudy (mathteacher1729):

And then subtracted one from the other: x = 6 - 6y - x =-7 + 6y ( I multiplied everything in this row by -1) -------------- 0 = -1 + 0 But 0 does not equal -1. Also, we can put the equations in y = mx + b form and see that the slopes (m) are equal, but the intercepts are not. That means we have two parallel lines, and parallel lines never intersect, so no solution. :)